A chemist must dilute 31.8mL of 4.67M aqueous sodium thiosulfate Na2S2O3 solution until the concentration falls to 3.00M . He'll do this by adding distilled water to the solution until it reaches a certain final volume. Calculate this final volume, in milliliters. Round your answer to 3 significant digits.
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 4.67 M

M2 = 3 M

V1 = 31.8 mL

use:

M1*V1 = M2*V2

V2 = (M1 * V1) / M2

V2 = (4.67*31.8)/3

V2 = 49.5mL

Answer: 49.5 mL